OpenStudy (anonymous):

If a loaded truck that can accelerate at 1m/s2 loses its load and has three fourths of its original mass, what acceleration can it obtain from the same driving force?

OpenStudy (anonymous):

Use \[\Sigma F = ma\] twice for the loaded truck (m1 a1) and unloaded truck (m2 a2) where a1 = 1 m/s^2 and m2 = (3/4)m1

OpenStudy (anonymous):

So in decimal form it is 0.75 m per second squared?

OpenStudy (compphysgeek):

the force accelerating your truck is constant. Now the mass is less, so the acceleration must be bigger. a = F/m. m2 is 3/4 of m1. so a2 is 4/3 of a1

OpenStudy (anonymous):

Ok I get it thanks, helped a lot